Diffusion Morning Report 10/20/04

tberg tberg at snf.stanford.edu
Wed Oct 20 06:46:39 PDT 2004

Good Morning All,
    Looking OK this morning. Nitride and Poly under observation. WB 
silicide is in the process of moving. Epi parts should be here this week.
Gas vault fans /sensor failed last night and all gases in the vaults 
went off-back on now.  Please check runs from 12:00am until 6:00am this 
morning for proper thicknesses.  TEOS2 and Thermco poly testing.  
Everything else seems ok for now.

. 	Date: 	10/20/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	Up 	  	12.33/30
Tylan 9-Poly 	yellow 	observation 	11.98//30
Tylan 10-Nitride 	yellow 	possible particle issues 	3.73/30
Tylan 11-BPSG-LTO 	Up 	  	23.07/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	7.46/30
Doped Poly 	GONE 	  	 
TEOS 	up 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	Testing 	 
Thermco Nitride 	Down 	Waiting 	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
SiH4 for furnaces 	  	Auto changeover working well 	B on line 510 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20041020/1337433f/attachment.html>

More information about the diffusion mailing list